Help me for the answer
3x+y=1
2x-y=-6

Answer by Alan3354(69443)

3x+y=1
2x-y=-6 Add the 2 equations
-------
5x = -5
x = -1
-------------
3x+y = 1
-3+y = 1
y = 4
--------------